OpenStudy (anonymous):

an automobile accelerates from zero to 30m/s in 6s. The wheels have a diameter of .4m. What is the angular acceleration of each wheel?

OpenStudy (anonymous):

\[a= \frac{\Delta v}{\Delta t}\] \[\alpha = \frac{\Delta \omega}{\Delta t}\] \[\omega = \frac{v}{r}\]

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

so 5/.4=12.5?

OpenStudy (anonymous):

oh, I'm so sorry. It didn't tell me that you had responded :( That looks right - remember the units though!!! ^_^ \[\frac{5 m/s^2}{.4m}=12.5 rad/s^2\]

OpenStudy (anonymous):

ooops oooops sorry sorry

OpenStudy (anonymous):

.4m is the diameter, and you want to use the radius = .2m

OpenStudy (anonymous):

ah so 5/.2= 25 rad/s^2

OpenStudy (anonymous):

yah ^_^

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

very welcome. Holler if you need anything else ^_^